OpenStudy (anonymous):

Can I multiply this: (3^n)*3 ?

OpenStudy (anonymous):

\[3^{n}\times3\]

OpenStudy (anonymous):

3^(n+1)

OpenStudy (anonymous):

\[3^{n+1}\]

OpenStudy (anonymous):

thanks, I was having this doubt with this problem, this is part of a series problem

OpenStudy (anonymous):

yea 3^{n+1} is the answer after u multiply

OpenStudy (anonymous):

Think about it this way: \[3^n \times 3 = \underbrace{\left(\underbrace{3 \times 3 \times \cdots \times 3}_{n \text{ times}}\right) \times \left(3\right)}_{n+1\text{ times}} = 3^{n+1}\]